# Stability of traveling waves for doubly nonlinear equations

### v2

@@ 1 Introduction and Main Theorem
In this work, we examine nonnegative solutions of (1) that are close to a traveling wave solution in order to get insights in the qualitative behavior of general solutions. As we are particularly interested in the behavior of the free boundary, considering traveling wave solutions appears to be natural. Indeed, traveling wave solutions exhibit the same boundary behavior as the so-called Barenblatt solutions, which are the radially symmetric self-similar solutions for (1) constructed first in [5]. These self-similar solutions do, similar to other parabolic problems, prescribe the large-time asymptotic behavior of any solution with finite mass, see [1, 2]. Moreover, it can be shown that any compactly supported solution of (1) has the same spreading behavior as the Barenblatt solutions [30]. Accordingly, localizing around an arbitrary boundary point of any solution would to leading order result in the problem on the half space we are considering here.
@@
Without loss of generality, we assume that the solution moves in the $n$th direction, that is $e=e_{n}$. In this case a traveling wave solution of (1) is defined by the profile
@@ 1.1 New coordinates and perturbation equation
In order to formulate our main result for the perturbation variable $w$ we have to introduce some notation. For its proof, we will make use of a fixed point argument. An idea, based Angenent’s trick [3] and an extension by Koch and Lamm [19] enables us to derive the analyticity statements. To perform the fixed point argument, we have to work with cleverly chosen Banach spaces. In [16], Kienzler introduced suitable Banach spaces $X$ and $Y$ (that we will define later), where a maximal regularity estimate of the form
@@
    For any initial datum $g\in X_{0}$, the linear initial value problem
@@
In contrast to the linear operator $\mathcal{L}_{\sigma}$, the nonlinearity $\mathcal{N}[w]$ in the perturbation equation (29) appears to be far more complicated compared to the porous medium equation case considered in [16]. However, a closer examination of the nonlinearity reveals enough (essentially quadratic) structure to establish assumption (A2) similar to [16, Lemma 2.3].
@@
Proof. We start by shortly addressing the analyticity of the nonlinearity $\mathcal{N}$. Since $\mathcal{N}$ is a quotient of algebraic functions, it is analytic away from its poles, which is located at $01+\partial_{n}w=$ 0. The proof of the stated estimates proceeds in the same way as the one of the corresponding lemma for the porous medium equation, see [16, Lemma 2.3]. We limit ourselves to show that the nonlinear terms in (29) satisfy the following bounds, from where the proof can be easily reproduced. That is, for $\left\lVert\nabla w\right\rVert_{L^{\infty}}\ll 1$ it holds that
@@
(59) $\left|\nabla\mathcal{N}[w]\right|\lesssim\left|\nabla w\right|\left|\nabla^{2}w\right|+z_{n}\left|\nabla w\right|\left|\nabla^{3}w\right|+z_{n}\left|\nabla^{2}w\right|^{2}.$
All occurring terms except the last term on the right-hand side of (59) can be easily controlled in the $X(q)$-norm. The remaining term can be interpolated via

### v3

@@ 1 Introduction and Main Theorem
In this work, we examine nonnegative solutions of (1) that are close to a traveling wave solution in order to get insights in the qualitative behavior of general solutions. As we are particularly interested in the behavior of the free boundary, considering traveling wave solutions appears to be natural. Indeed, traveling wave solutions exhibit the same boundary behavior as the so-called Barenblatt solutions, which are the radially symmetric self-similar solutions for (1) constructed first in [6]. These self-similar solutions do, similar to other parabolic problems, prescribe the large-time asymptotic behavior of any solution with finite mass, see [1, 2]. Moreover, it can be shown that any compactly supported solution of (1) has the same spreading behavior as the Barenblatt solutions [35]. Accordingly, localizing around an arbitrary boundary point of any solution would result to leading order in the problem on the half space we are considering here.
Traveling waves display the typical behavior of solutions also in the presence of reactions terms, as studied, for instance, in [16, 5, 14, 15, 12]. The profiles of these waves typically decay to zero at infinity, and consequently, these solutions lack a free boundary, which simplifies their analysis.
@@
Without loss of generality, we assume that the solution moves in the $n$th direction, that is $e=e_{n}$. In this case a traveling wave solution of (1) can be defined by the profile
@@ 1.1 New coordinates and perturbation equation
Here, we have used the notation $\nabla^{\prime}=(\partial_{1},\dots,\partial_{n-1})^{T}$ for the tangential gradient and $\Delta^{\prime}=\partial_{1}^{2}+\dots\partial_{n-1}^{2}$ for the tangential Laplacian.
@@
In order to formulate our main result for the perturbation variable $w$ we have to introduce some notation. For its proof, we will make use of a fixed point argument. An idea based Angenent’s trick [3] and an extension by Koch and Lamm [24] enables us to derive the analyticity statements. To perform the fixed point argument, we have to work with cleverly chosen Banach spaces. In [21], Kienzler introduced suitable Banach spaces $X$ and $Y$ (that we will define later), where a maximal regularity estimate of the form
@@
    For any initial datum $g\in X_{0}$ and $f\in Y$, the linear initial value problem
@@
In contrast to the linear operator $\mathcal{L}_{\sigma}$, the nonlinearity $\mathcal{N}[w]$ in the perturbation equation (29) appears to be far more complicated compared to the porous medium equation case considered in [21]. However, a closer examination of the nonlinearity reveals enough (essentially quadratic) structure to establish assumption (A2) similar to [21, Lemma 2.3].
@@
Proof. We start by shortly addressing the analyticity of the nonlinearity $\mathcal{N}$. Since $\mathcal{N}$ is a quotient of algebraic functions, it is analytic away from its poles, which is located at $01+\partial_{n}w=$ 0. The proof of the stated estimates proceeds in the same way as the one of the corresponding lemma for the porous medium equation, see [21, Lemma 2.3]. We limit ourselves to show that the nonlinear terms in (29) satisfy the following bounds, from where the proof can be easily reproduced. That is, for $\left\lVert\nabla w\right\rVert_{L^{\infty}}\ll 1$ it holds that
@@
(59) $\left|\nabla\mathcal{N}[w]\right|\lesssim\left|\nabla w\right|\left|\nabla^{2}w\right|+z_{n}\left|\nabla w\right|\left|\nabla^{3}w\right|+z_{n}\left|\nabla^{2}w\right|^{2},$
where $\nabla^{2}w$ and $\nabla^{3}w$ are the matrices of all second and third order derivatives, respectively. All occurring terms except the last term on the right-hand side of (59) can be easily controlled in the $X(q)$-norm. The remaining term can be interpolated via
@@
Acknowledgment
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 –390685587, Mathematics Münster: Dynamics–Geometry–Structure.